OpenStudy (anonymous):

enter the quadratic equation in standard form x^2-10x-6=2x+3

OpenStudy (lgbasallote):

just subtract both sides by (2x+3) what do you get?

OpenStudy (anonymous):

no :( explain

OpenStudy (lgbasallote):

hmm okay x^2 -10x -6 = 2x + 3 subtract both sides by 2x

OpenStudy (anonymous):

idk how to do that lol

OpenStudy (anonymous):

x^2-4x+3?

OpenStudy (lgbasallote):

x^2 - 10 x - 6 - 2x = 2x + 3 - 2x so this becomes x^2 - 10x - 2x - 6 = 3 got it?

OpenStudy (lgbasallote):

and no not what you said

OpenStudy (anonymous):

sorta not really

OpenStudy (lgbasallote):

what's 10-2?

OpenStudy (anonymous):

8?

OpenStudy (lgbasallote):

right. so what is 10x - 2x?

OpenStudy (anonymous):

8x?

OpenStudy (lgbasallote):

correct. now..what is -10x -2x?

OpenStudy (anonymous):

12x

OpenStudy (anonymous):

-8x!!!

OpenStudy (lgbasallote):

no..not -8x

OpenStudy (lgbasallote):

-10x - 2x is similar to 10x + 2x but with a negative sign

OpenStudy (anonymous):

so 12x?

OpenStudy (lgbasallote):

that's 10x + 2x.... what is 10x + 2x with a negaive sign?

OpenStudy (anonymous):

-12x

OpenStudy (lgbasallote):

right

OpenStudy (lgbasallote):

so now what is x^2 - 10x - 2x - 6 = 3

OpenStudy (anonymous):

x^2-12x-9=0?!!!

OpenStudy (lgbasallote):

you got it